OpenStudy (anonymous):

How do u factor 2x^3+3x^2-8x-12

OpenStudy (anonymous):

anyone have an answer?

OpenStudy (anonymous):

or to (16n^4)-1

OpenStudy (anonymous):

factored as well

OpenStudy (zzr0ck3r):

x^2(2x+3)-4(2x+3) you got the rest?

OpenStudy (anonymous):

yep, thanks so much :) btw, would you happen to know how to factor the second problem?

OpenStudy (anonymous):

OH. okay. I hope this isn't wrong, because I get things wrong....a lot.. But use factor by grouping. First group the 4 into 2 pairs, such as: (2x^3+3x^2) + (-8x-12) Factor these x^2(2x+3) -4(2x+3) Put together. (x^2 -4) (2x+3) You can still factor the first one, so: (x-2)(x+2)(2x+3)

OpenStudy (anonymous):

that looks great, thanks to both of you

OpenStudy (anonymous):

anyone happen to know (16n^4)-1?

OpenStudy (anonymous):

factored

OpenStudy (zzr0ck3r):

((4n)^2-1^2) remember (a^2-b^2) = (a-b)(a+b)

OpenStudy (zzr0ck3r):

so first step (((4n)^2)^2-1^2) = ((4n)^2-1)((4n)^2+1) second sterp hint you can do it again with (4n)^2-1

OpenStudy (anonymous):

thanks for the assistance, i deeply appreciate it